OpenStudy (anonymous):

Evaluate the following expression for x = -8 and y = 5 5y – 3(5 – x) + 7x + 4(3x – 2y)

OpenStudy (anonymous):

replace all of the x's with -8 and the y's with 5

OpenStudy (anonymous):

\[5(5)-3(5--8)+7(-8)+4(3(-8)-2(5))\]

OpenStudy (anonymous):

now use order of operations

OpenStudy (vishweshshrimali5):

5*5 -3(5-(-8) + 7(-8) +3(3(-8) -2(5)) = 25 -3(13) -56 + 3(-24 -10) = 25 - 39 - 56 + 3(-34) = 25 - 39 - 56 - 102 = -172

OpenStudy (anonymous):

thanks vish im done for today

OpenStudy (vishweshshrimali5):

its ok